i am trying to get the equation for a mixture problem.
You have 7 ounces of 30% glucose solution. How much 30% glucose solution and how much water needs to be used to make 7 ounces of 20% solution?
my equation i tried is
.30(7) + x = .20(7-x)
is this the right equation to start the problem?
Thank you for the help

Answer by stanbon(75887)

You have 7 ounces of 30% glucose solution. How much 30% glucose solution and how much water needs to be used to make 7 ounces of 20% solution?
----------------
Equation:
sugar + sugar = sugar
0.30*7 + 0*x = 0.20(7+x)
----
Multiply thru by 100 to get:
30*7 + 0 = 20*7 + 20x
20x = 10*7
x = 3.5 ounces (amt. of water that should be added)
=================================================
Proof that this is the correct answer:
0.3*7 = 2.1 oz of sugar
0.2(7+3.5) = 0.2(10.5) = 2.1 oz of sugar
=====================
Cheers,
Stan H.
=====================